A generator is connected to a resistor and a 0.029-H inductor in series. The rms voltage across the generator is 7.6 V. When the generator frequency is set to 134 Hz, the rms voltage across the inductor is 2.7 V. Determine the resistance of the resistor in this circuit. Ω
Answer #1

The reactance of the inductor = 2πf*L = 2π*134*0.029 = 24.41Ω

So the current through the inductor = V/XC = 2.7/24.41 = 0.1106A

We have V (generator) = sqrt(Vr^2 + VL^2) => Vr = sqrt(V^2 - vL^2) = sqrt(7.6^2 - 2.7^2) = 7.10V

So R = V/I = 7.10/0.1106 = 64.23Ω = 64Ω